# Mitigating photon loss in linear optical quantum circuits: classical postprocessing methods outperforming postselection

### v2

@@ 8.2 Extrapolation
Note that the ideal signal magnitude decays proportionally with (${{m-n+k}\choose k}^{-1}\sim m^{-k}$ in eqn. 28, which intuitively motivates the choice of an exponential model function to reflect this decay behaviour. In the next sections, we provide numerical evidence indicating that there exists a non-trivial sampling regime where exponential extrapolation outperforms postselection. We also note that in the numerical simulations, extrapolation using an exponential model function consistently outperforms linear extrapolation (see Fig. 3 (c) and (d)). This may be a consequence of the exponential model function better reflecting the ideal signal decay behaviour.
9 Numerical computation of regime of mitigation usefulness for example experiment
To demonstrate how the performance guarantees may be related to a real experiment, we now show how the performance guarantees for linear solving recycling mitigation may be used to compute a regime of usefulness where the mitigation outperforms postselection. Let the example experiment consist of a computation with size $m=100$ modes, $n=10$ photons, that is run on noisy hardware for which the uniform probability of photon loss is $\eta=0.8$, and for the linear solving method let $k=1$. We now compute a regime of mitigation usefulness for an arbitrary probability from an arbitrary distribution. The performance guarantee for linear solving recycling mitigation of a single probability is  [^4] Note that the statistical errors in eqn. 36 and Thm. 10 differ by a numerator ($\sqrt{m\choose n}$, this is because in our proof of Thm. 10 we divided the sample size by (${m\choose n}$ to ensure we are computing each recycled probability from an independent sample group (see appendix 18). Here since we are interested in a single probability, we drop this division step.
$\underbrace{O\Bigg{(}\sqrt{\frac{1}{{n\choose k}(1-\eta)^{n-k}\eta^{k}N_{tot}}}\Bigg{)}}_{\text{Mitigation statistical error}}+\underbrace{\bigg{(}{m-n+k\choose k}-1\bigg{)}O\bigg{(}\frac{1}{\text{poly}(m)}\bigg{)}}_{\text{Mitigation bias error}}\leq\underbrace{O\Bigg{(}\sqrt{\frac{1}{(1-\eta)^{n}N_{tot}}}\Bigg{)}}_{\text{Postselection statistical error}}.$ (36)
We need to replace the big O notation with prefactors to allow computation of a bound. So first the polynomial interference term bias error is set to $\epsilon_{bias,s^{n}_{l}}=\frac{1}{m^{2k}}$. And next the confidence for both mitigated and postselected statistical error bounds is set to $1-\delta$, $\delta={m}^{-n}$. Now from Hoeffding’s inequality [26], with $N_{est,k}\approx{n\choose k}\eta^{k}(1-\eta)^{n-k}N_{tot}$, we have
$2e^{-2N_{est,k}\epsilon_{\text{hoeff}}^{2}}\geq{m}^{-n}.$ (37)
Taking the $\log$ of both sides
$-2N_{est,k}\epsilon_{\text{hoeff}}^{2}\geq-\log(2)-n\log(m),$ (38)
and rearranging this becomes
$\epsilon_{\text{hoeff}}\geq\sqrt{\frac{\log(2)+n\log(m)}{2N_{est,k}}}.$ (39)
This leads to a polylogarithmic prefactor term for the statistical error bounds for the mitigated and postselected outputs of $\sqrt{2^{-1}(\log 2+n\log(m))}$. We then get the following condition for an arbitrary probability in the mitigated distribution to be more accurate than the corresponding probability in the postselected distribution
$3\sqrt{\frac{\big{(}\log 2+n\log(m)\big{)}}{2{n\choose k}(1-\eta)^{n-k}\eta^{k}N_{tot}}}+3\bigg{(}{m-n+k\choose k}-1\bigg{)}\frac{1}{m^{2k}}\leq\sqrt{\frac{\big{(}\log 2+n\log(m)\big{)}}{2(1-\eta)^{n}N_{tot}}}.$ (40)
This can then be rearranged as
$N_{tot}\leq\frac{\big{(}\log 2+n\log(m)\big{)}m^{4k}}{2\big{(}{m-n+k\choose k}-1\big{)}^{2}}\bigg{(}\frac{1}{9(1-\eta)^{n}}-\frac{1}{{n\choose k}(1-\eta)^{n-k}\eta^{k}}\bigg{)}.$ (41)
Substituting in the parameters of the example experiment we get that, with exponentially high confidence in terms of $n$, linear solving recycling mitigation should outperform postselection for at least up to $2.42\times 10^{11}$ samples. The upper bound on the usefulness sampling regime for these experimental parameters, only with photon number varying in the range $n\in[5,10]$, is plotted in Fig. 2.
@@ 10 Numerical simulations
Figure 2: The upper bounds on the total number of samples for which linear solving recycling mitigation will outperform postselection plotted against photon number. These values were computed from the performance guarantees for the parameter values $m=100$, $\eta=0.8$, $k=1$, and with $n\in[5,10]$. A bias error of $\epsilon_{bias,s^{n}_{l}}=m^{-2k}$ and confidence for the statistical error bounds of $m^{-n}$, which gives rise to a logarithmic prefactor term for the statistical error bounds of $\sqrt{2^{-1}\log 2+n\log(m)}$. This displays the exponential dependence on the photon number of the upper bound on the total number of samples for which the techniques will outperform postselection.
Figure 3: A numerical performance comparison of linear solving and postselection for random unitary circuits with $m=20$ modes and $n=4$ photons. (a) The KL divergence of the outputs of linear solving recycling mitigation and postselection from the ideal outputs for a uniform loss parameter of $\eta=0.8$ against total sample number. (b) The KL divergence for random unitary circuits with a total number of samples of $N_{tot}=1\times 10^{5}$, and with the uniform loss parameter in the range $\eta\in[0.5,0.9]$. (c) KL divergence for random unitary circuits with 20 modes, 4 photons, a uniform loss parameter of 0.8 against total sample number. (d) KL divergence for random unitary circuits with a total number of samples of $N_{tot}=1\times 10^{5}$, and with the uniform loss parameter in the range $\eta\in[0.5,0.9]$.
In the numerical simulations recycling mitigation is used to mitigate the effects of uniform photon loss error on the output of an otherwise ideal simulation of a linear optical quantum circuit. In the simulations, random matrices are chosen for each experiment which are decomposed and implemented with a linear interferometer, and a uniform photon loss model is applied with a defined loss parameter. The simulations were performed using Perceval [20], a pythonic framework for the simulation of photonic quantum circuits. Uniform photon loss channels commute with the interferometer, and so all loss channels, including loss due to imperfect sources and measurement, can be propagated to the end of the circuit. And so the effects of loss may be modelled by an ideal photon source, an ideal interferometer $M$, and a combined photon loss channel acting immediately before an ideal measurement operation. This means that rather than an output photon from the circuit incident on a detector being in the state |$\ket{1}\bra{1}$ |, it is instead
@@
Figure 4: The upper bound for the total number of samples for which linear solving recycling mitigation will outperform postselection according to Thm. 10. The experimental parameters used were $m=20$, $n=4$ and $\eta\in[0.75,0.975]$. The data point for $\eta=0.8$ may be seen to be in agreement with the data in Fig. 3 (a). Here for an arbitrary single probability the sample upper bound for which the worst case error for linear solving recycling mitigation is lower than that for postselection for $\eta=0.8$ is $\approx 1.1\times 10^{5}$. Whereas in Fig. 3 (a) where the full distribution is mitigated it is $\approx 1.1\times 10^{6}$ samples.
For all the experiments in Fig. 3, parameter settings of 20 modes and 4 photons were used. Fig. 3 (a) plots the performance of linear solving recycling mitigation, both with and without using the dependency factor, and postselection against the total number of samples used for a fixed uniform loss parameter of $\eta=0.8$. Linear solving outperformed postselection up to $\approx 1.1\times 10^{6}$ samples, while linear solving with dependency outperformed postselection up to $\approx 3.1\times 10^{6}$ samples. Fig. 3 (b) plots the performance of the linear solving methods and postselection against changing loss parameter for a fixed number of samples of $1\times 10^{5}$. Linear solving outperformed postselection for loss above $\approx 0.6$, and linear solving with the dependency factor above loss of $\approx 0.5$. Fig. 3 (c) plots the performance of linear extrapolation, exponential extrapolation and postselection against the total number of samples used for a fixed uniform loss parameter of $\eta=0.8$. Linear extrapolation outperformed postselection up to $\approx 1.7\times 10^{6}$ samples, while exponential extrapolation does so up to $\approx 3.0\times 10^{6}$ samples. Fig. 3 (d) plots the performance of the extrapolation methods and postselection against changing loss parameter for a fixed number of samples of $1\times 10^{5}$. Linear extrapolation outperformed postselection for loss above $\approx 0.5$, and linear solving with the dependency factor above loss of $\approx 0.5$. In current quantum linear optical devices losses including source, interferometer and measurement are commonly above $50$% [18]. Therefore, the evidence from these simulations indicates that recycling mitigation may be usefully applied to the current generation of linear optical devices. Also, in the previous section an upper bound on the number of samples was computed for a particular experiment. Fig. 4 was plotted in order to compare the analytical expression in eqn. 41 to the numerics in 3 (a). In Fig. 4, for experimental parameters $m=20$, $n=4$ and $\eta\in[0.75,0.975]$, the sample upper bound for which the worst-case error for linear solving recycling mitigation for an arbitrary single probability is lower than that for postselection for $\eta=0.8$ is $\approx 1.1\times 10^{5}$. Whereas in Fig. 3 (a) where the full distribution is mitigated it is $\approx 1.1\times 10^{6}$ samples. And so the mitigation performance prediction based on eqn. 41 is in agreement with the numerics. Compared with the numerics, eqn. 41 estimates a lower sample threshold up to which the mitigation will outperform postselection. The most likely explanation for this is that the bias error is lower in the numerical experiments than is assumed for the analytics.
@@
Theorems 3 and 5 in section 7 provide statistical upper bounds on the deviation of the interference terms from $p_{unif}$. We have shown that with confidence $\approx 1-O^{*}(m^{-n})$ these upper bounds are within a range $[-O\big{(}\frac{1}{\mathsf{poly}(m)}\big{)},O\big{(}\frac{1}{\mathsf{poly}(m)}\big{)}]$ around $p_{unif}$. We conjecture that these bounds could be considerably tightened. In support of this we ran numerical experiments for computations of size $m=16$ modes and $n=4$ photons, computing the average of $|I_{s^{n}_{l},k}-p_{unif}|$ over all bit strings $\{s^{n}_{l}\}_{l}$. This computation was repeated for 20 randomly selected linear optical intereferometers $U$. This data is plotted in Fig. 5. It may be observed that for all unitaries, and both when $k=1$ and when $k=2$, the computed values were of the order of (${m\choose n}^{-1}$, and so were in fact exponentially small in terms of the size of the computation. This indicates that the analytical performance guarantees given in section 8 may under-estimate the size of the sample number up to which recycling mitigation will outperform postselection. Finally, note that $|I_{s^{n}_{l},k}-p_{unif}|$ seems to be generally decreasing with increasing $k$. However, (${m-n+k\choose k}|I_{s^{n}_{l},k}-p_{unif}|$ actually increases with increasing $k$, as can be verified by a direct calculation using data from Figure 5. Ultimately, since (${m-n+k\choose k}|I_{s^{n}_{l},k}-p_{unif}|$ is directly related to the bias error incurred when computing the mitigated probabilities from the recycled probabilities (see section 8), increasing values of $k$ in the recycling mitigation will lead to worse performances, as predicted by Lemma 1. We conjecture that (${m-n+k\choose k}|I_{s^{n}_{l},k}-p_{unif}|$ increases monotonically with increasing $k$ (see section 14).
@@
Figure 5: The mean absolute value of the distance of the interference terms from the uniform probability were computed for 20 randomly selected unitaries. For all unitaries, and for $k=1$ and $k=2$, the magnitude of the computed values were observed to be exponentially small in terms of $m$ and $n$ (being of the order of (${m\choose n}^{-1}$). This indicates that it may be possible to derive tighter analytical bounds than those stated in thm. 3 and thm. 5, and that mitigation performance may be better than indicated by the performance guarantee statements given in section 8.
@@ 11 Evidence that zero noise extrapolation (ZNE) techniques present no advantage over postselection
[…]
@@
Thm. 13, proven in appendix 21, is strong evidence that techniques based on Richardson extrapolation offer no advantage over postselection. Our main technical contribution in proving Thm. 13 is to link determining the error of ZNE methods to computing the norm of the inverse of a Vandermonde matrix whose entries are determined by $\eta$, we then use existing results to upper bound this norm [65]. Furthermore, in appendix 21 we also provide numerical evidence to support this conclusion.
12 On normalising the mitigated distribution
The postprocessing techniques we have presented thus far take as input a set of recycled probabilities $\{p_{R}(s^{n}_{l})\}_{l}$ and output a set of mitigated values $\{p_{mit}(s^{n}_{l})\}_{l}$, or a subset of these values. We have made the distinction between values and probabilities, as the outputs $\{p_{mit}(s^{n}_{l})\}_{l}$ satisfy $p_{mit}(s^{n}_{l})>0$ but are not normalised in general. That is, $\sum_{l}p_{mit}(s^{n}_{l})=N$ with no guarantee that $N=1$. Analytical and numerical calculations performed in previous sections have shown that, with high probability,
$||\vec{p}_{mit}-\vec{p}_{id}||_{1}\leq||\vec{p}_{post}-\vec{p}_{id}||_{1},$
and $|p_{mit}(s^{n}_{l})-p(s^{n}_{l})|\leq|p_{post}(s^{n}_{l})-p(s^{n}_{l})|$ as long as we are in a sampling regime defined by eqn. (23). $\vec{p}_{mit}$, $\vec{p}_{id}$, and $\vec{p}_{post}$, are vectors containing respectively the mitigated values, the ideal $n$–photon probabilities, and the probabilities obtained by postselection, and ||.$||.||_{1}$ is the usual $l_{1}$ norm, $||\vec{v}||_{1}:=\sum_{i}|v_{i}|$.
In practice, one is usually interested in computing the expectation value of some observable $\mathsf{O}$, defined as $\langle\mathsf{O}\rangle=\sum_{i}p_{i}w_{i}$, where $w_{i}$ are some weights, and $\{p_{i}\}$ a subset of probabilities of the quantum circuit. In this case we can replace the $p_{i}$’s by the unnormalised mitigated values and obtain guarantees similar to those stated above. However, if we are interested in mitigating the entire distribution, then we would need to normalise $\vec{p}_{mit}$. The easiest way to do this, and what we do for our numerical simulations, is to define, $\vec{p}_{mit,nor}:=\frac{1}{N}\vec{p}_{mit}$. One can easily check that $||\vec{p}_{mit,nor}||_{1}=1$, and therefore that it is a vector of normalised probabilities.
We will show that the normalised probabilities $\vec{p}_{mit,nor}$ exhibit similar guarantees to the unnormalised mitigated values $\vec{p}_{mit}$, albeit in a sampling regime which is slightly more restrictive than eqn. (23). We believe, however, that this restriction is just an artifact of the proof technique. Let $\vec{q}_{m}$ be a vector of positive values which is a closest (in $l_{1}$-norm) normalised vector to $\vec{p}_{mit}.$ More precisely, $\vec{q}_{m}$ satisfies
$||\vec{q}_{m}-\vec{p}_{mit}||_{1}=\mathsf{min}_{\vec{q}\big{|}||\vec{q}||_{1}=1}||\vec{q}-\vec{p}_{mit}||_{1}.$
An immediate observation, by definition, is that
$||\vec{q}_{m}-\vec{p}_{mit}||_{1}\leq||\vec{p}_{id}-\vec{p}_{mit}||_{1}.$
We now prove

### v3

@@ 4.4.2 Extrapolation
Note that the ideal signal magnitude decays proportionally with ${{m-n+k}\choose k}^{-1}\sim m^{-k}$ in eqn. 31, which intuitively motivates the choice of an exponential model function to reflect this decay behaviour. In the next sections, we provide numerical evidence indicating that there exists a non-trivial sampling regime where exponential extrapolation outperforms postselection. We also note that in the numerical simulations, extrapolation using an exponential model function consistently outperforms linear extrapolation (see Fig. 3 (c) and (d)). This may be a consequence of the exponential model function better reflecting the ideal signal decay behaviour.
As a final remark we comment on how one can estimate $D_{k}$ in practice. In [2], it was shown that $D_{0}$ is lower bounded by a constant. The results of numerical simulations plotted in Fig. 2 highlight that the same appears to hold for $D_{k}$ when $k\in O(1)$. Intuitively, this makes sense as $D_{k}$, similar to $D_{0}$, are expected to contain probabilities that are hard to simulate classically (see Section 6) , and are thus far from the uniform distribution. In practice, this means that in order to estimate $D_{k}$ to a good ($\epsilon\in O(1)$) precision, one only needs to compute an $O(1)$ fraction of probabilities $p_{R}(s^{n}_{l})$, with $s^{n}_{l}$ picked uniformly randomly, and compute the mean of $|p_{R}(s^{n}_{l})-p_{unif}|$. This gives a high-confidence estimate of $D_{k}$ by Hoeffding’s inequality. Consequently, this means that estimating the $D_{k}$ to be used in extrapolation recycling mitigation does not require significantly more statistics than what is needed in linear solving.
Figure 2: The results of numerical simulations where $D_{k}$ for $k\in\{1,2\}$ is plotted for $30$ random unitaries with $m=20$ for (a) $n=3$, (b) $n=4$, (c) $n=5$, (d) $n=6$, and (e) $n=7$. From these results, as with [2] where it was shown that $D_{0}$ is lower bounded by a constant, it appears this also holds for $D_{k}$ when $k\in O(1)$.
4.5 Normalising the mitigated distribution
The postprocessing techniques we have presented take as input a set of recycled probabilities $\{p_{R}(s^{n}_{l})\}_{l}$ and output a set of mitigated values $\{p_{mit}(s^{n}_{l})\}_{l}$, or a subset of these values. We have made the distinction between values and probabilities, as the outputs $\{p_{mit}(s^{n}_{l})\}_{l}$ satisfy $p_{mit}(s^{n}_{l})>0$ but are not normalised in general. That is, $\sum_{l}p_{mit}(s^{n}_{l})=N$ with no guarantee that $N=1$. Analytical and numerical calculations performed in previous sections have shown that, with high probability,
[…]
and $|p_{mit}(s^{n}_{l})-p(s^{n}_{l})|\leq|p_{post}(s^{n}_{l})-p(s^{n}_{l})|$ up to a certain number of samples. $\vec{p}_{mit}$, $\vec{p}_{id}$, and $\vec{p}_{post}$, are vectors containing respectively the mitigated values, the ideal nn–photon probabilities, and the probabilities obtained by postselection, and $||.||_{1}$ is the usual $l_{1}$ norm, $\vec{v}||_{1}:=\sum_{i}|v_{i}|$.
In practice, one is usually interested in computing the expectation value of some observable $\mathsf{O}$, defined as $\langle\mathsf{O}\rangle=\sum_{i}p_{i}w_{i}$, where $w_{i}$ are some weights, and $\{p_{i}\}$ a subset of probabilities of the quantum circuit. In this case we can replace the $p_{i}$’s by the unnormalised mitigated values and obtain guarantees similar to those stated above. However, if we are interested in mitigating the entire distribution, then we would need to normalise $\vec{p}_{mit}$. The easiest way to do this, and what we do for our numerical simulations, is to define, $\vec{p}_{mit,nor}:=\frac{1}{N}\vec{p}_{mit}$. One can easily check that $\vec{p}_{mit,nor}||_{1}=1$, and therefore that it is a vector of normalised probabilities.
We will show that such a normalisation provably produces a distribution $\vec{p}_{mit,nor}$ which is closest possible to $\vec{p}_{mit}$ in $l_{1}$-norm.
Let $\vec{q}_{m}$ be a vector of positive values which is a closest (in $l_{1}$-norm) normalised vector to $\vec{p}_{mit}.$ More precisely, $\vec{q}_{m}$ satisfies
[…]
An immediate observation, by definition, is that
[…]
We now prove
Lemma 15.
A valid choice of $\vec{q}_{m}$ is $\vec{q}_{m}=\vec{p}_{mit,nor}$.
Proof.
[…]
@@ 5 Numerical simulations
Figure 3: A numerical performance comparison of the different methods of recycling mitigation and postselection for random unitary circuits with $m=20$ modes and $n=4$ photons. (a) For a uniform loss parameter of $\eta=0.8$, the KL divergence from the ideal output distribution for linear extrapolation, exponential extrapolation, linear solving and linear solving with dependency versions of recycling mitigation and postselection is plotted against total sample number. (b) For a total number of samples of $N_{tot}=1\times 10^{5}$ and a uniform loss parameter in the range $\eta\in[0.5,0.9]$, the KL divergence from the ideal output distribution of linear extrapolation, exponential extrapolation, linear solving and linear solving with dependency versions of recycling mitigation and postselection is plotted against total sample number.
In the numerical simulations recycling mitigation is used to mitigate the effects of uniform photon loss error on the output of an otherwise ideal simulation of a linear optical quantum circuit. In the simulations, random matrices are chosen for each experiment which are decomposed and implemented with a linear interferometer, and a uniform photon loss model is applied with a defined loss parameter. The simulations were performed using Perceval [24], a pythonic framework for the simulation of photonic quantum circuits. Uniform photon loss channels commute with the interferometer, and so all loss channels, including loss due to imperfect sources and measurement, can be propagated to the end of the circuit. And so the effects of loss may be modelled by an ideal photon source, an ideal interferometer $M$, and a combined photon loss channel acting immediately before an ideal measurement operation. This means that rather than an output photon from the circuit incident on a detector being in the state $|\ket{1}\bra{1}$, it is instead
@@
For all the experiments in Fig. 3, parameter settings of 20 modes and 4 photons were used. Fig. 3 (a) plots the performance of linear solving recycling mitigation, both with and without using the dependency factor, and postselection against the total number of samples used for a fixed uniform loss parameter of $\eta=0.8$. Linear solving outperformed postselection up to $\approx 1.1\times 10^{6}$ samples, while linear solving with dependency outperformed postselection up to $\approx 3.1\times 10^{6}$ samples. Note that ${20\choose 4}^{2}\approx 23\times 10^{6}$, and thus recycling mitigation is outperforming postselection for sample sizes of the order of ${m\choose n}^{2}={20\choose 4}^{2}$. Fig. 3 (b) plots the performance of the linear solving methods and postselection against changing loss parameter for a fixed number of samples of $1\times 10^{5}$. Linear solving outperformed postselection for loss above $\approx 0.6$, and linear solving with the dependency factor above loss of $\approx 0.5$. Fig. 3 (c) plots the performance of linear extrapolation, exponential extrapolation and postselection against the total number of samples used for a fixed uniform loss parameter of $\eta=0.8$. Linear extrapolation outperformed postselection up to $\approx 1.7\times 10^{6}$ samples, while exponential extrapolation does so up to $\approx 3.0\times 10^{6}$ samples. Fig. 3 (d) plots the performance of the extrapolation methods and postselection against changing loss parameter for a fixed number of samples of $1\times 10^{5}$. Linear extrapolation outperformed postselection for loss above $\approx 0.5$, and linear solving with the dependency factor above loss of $\approx 0.5$. These results indicate the exponential function is a better model for the physical behaviour of the decay than the linear function, we believe this may be explained in terms of the prefactor of the ideal output probability in eqn. 31 decreasing exponentially (specifically $\propto(m-n+k)^{-k}\big)$ with increasing $k$. In current quantum linear optical devices losses including source, interferometer and measurement are commonly above 5050% [40]. Therefore, the evidence from these simulations indicates that recycling mitigation may be usefully applied to the current generation of linear optical devices. Also, in the previous section an upper bound on the number of samples was computed for a particular experiment.
@@
Theorems 8 and 10 in section 4.3 provide statistical upper bounds on the deviation of the interference terms from $p_{unif}$. An inspection of these bounds shows that this deviation is with high confidence $\Omega(\sqrt{p_{unif}})$.
We conjecture that this bound could be considerably tightened. In support of this we ran numerical experiments for computations of size $m=16$ modes and $n=4$ photons, computing the average of $|I_{s^{n}_{l},k}-p_{unif}|$ over all bit strings $\{s^{n}_{l}\}_{l}$. This computation was repeated for 20 randomly selected linear optical intereferometers $U$. This data is plotted in Fig. 4. As can be seen from the figure it seems that the analytically computed bounds on the bias error are loose. Indeed, it seems that $|I_{s^{n}_{l},k}-p_{unif}|\in o(p_{unif})$, in line with conjecture 3.
Finally, note that $|I_{s^{n}_{l},k}-p_{unif}|$ seems to be generally decreasing with increasing $k$. However, ${m-n+k\choose k}|I_{s^{n}_{l},k}-p_{unif}|$ actually increases with increasing $k$, as can be verified by a direct calculation using data from Figure 4. Ultimately, since ${m-n+k\choose k}|I_{s^{n}_{l},k}-p_{unif}|$ is directly related to the bias error incurred when computing the mitigated probabilities from the recycled probabilities (see appendix 12), increasing values of $k$ in the recycling mitigation will lead to worse performances, as predicted by Lemma 6.
@@
Figure 4: The mean absolute value of the distance of the interference terms from the uniform probability were computed for 20 randomly selected unitaries. For all unitaries, and for $k=1$ and $k=2$, the magnitude of the computed values were observed to be exponentially small in terms of $m$ and $n$ (being of the order of ${m\choose n}^{-1}$). This indicates that it may be possible to derive tighter analytical bounds than those stated in thm. 8 and thm. 10.
6 Properties of the mitigated probabilities in the absence of statistical error
Figure 5: The numerical simulations that produced these plots involved exact computation of the mitigated probabilities and so only the bias error is present. The magnitude of the average error per output bit string is compared with that of the uniform probability. For all chosen parameters the average error per bit string is below the uniform probability. (a) The average error per bit string plotted for a range of modes $[16,30]$ for a fixed number of input photons $n=5$. (b) The average error per bit string plotted for a range of input photons $[3,9]$ for a fixed number of modes photons $m=20$.
As seen previously, recycling mitigation is a biased error mitigation. Given a $U$ and a bit string $s$, this means that in the limit of infinite samples the mitigated probability $p(s)$ is related to the ideal probability $p_{id}(s)$ by
$|p(s)-p_{id}(s)|=\epsilon_{{mitbias},s}.$
Numerical simulations in Fig. 5 using the exponential extrapolation recycling mitigation technique with values of $k$ independent of $m,n$ show that $E_{s}(\epsilon_{{mitbias},s})<p_{unif}$ both in the case of fixed $m$ and varying $n$, as well as fixed $n$ and varying $m$. Although the plots in Fig. 5 are for a fixed $U$, we observe similar curves when testing with different values of $U$. Our numerical simulations suggest that Conjecture 3 is true. We will now assume that this conjecture is true and derive from it a condition on the mitigated probabilities.
Let $m=\Omega(n^{5})$, then $p_{id}(s)\approx\frac{|\mathsf{Per}(X_{s})|^{2}}{m^{n}}$ with $X_{s}$ an i.i.d Gaussian matrix [1], also $p_{unif}\approx\frac{n!}{m^{n}}$. Thus, Conjecture 3 implies, by using a Markov inequality with $\alpha>1$,
$Pr_{s}(|m^{n}p(s)-|\mathsf{Per}(X_{s})|^{2}|\leq\delta(n)\alpha n!)\geq 1-\frac{1}{\alpha}.$
Choosing $\alpha$ such that $\alpha\delta(n):=\gamma(n)<1$, we have that
$Pr_{s}(|m^{n}p(s)-|\mathsf{Per}(X_{s})|^{2}|\leq\gamma(n)n!)\geq 1-\frac{1}{\alpha}.$
Since this equation holds with high probability $p$ over the choice of $U$, then we can say that
$Pr_{X\in\mathcal{G}^{n\times n}}(|m^{n}p(s)-|\mathsf{Per}(X)|^{2}|\leq\gamma(n)n!)\geq(1-\frac{1}{\alpha})p,$ (38)
where $\mathcal{G}^{n\times n}$ is the ensemble of $n\times n$ Gaussian matrices.
[…]
In conclusion, despite recycling mitigation being a biased error mitigation technique, we have presented numerical evidence to argue that the bias errors are small enough to render the mitigated probabilities hard to compute classically.
@@ 7 Evidence that zero noise extrapolation (ZNE) techniques present no advantage over postselection
In this section, we provide strong evidence that techniques based on ZNE when applied to mitigating photon loss in DVLOQC in general provide no advantage over postselection.
Suppose we are interested in computing a specific marginal probability $p(n_{1}...n_{l}|n)$ of observing $n_{i}$ photons in mode $i\in\{1,\dots,l\}$, with $l\leq m$, and $\sum_{i=1,\dots,l}n_{i}=c$, with $c\leq n$. The notation $|n$ indicates that we are computing the ideal marginal probability, when no photon is lost. Let $p(n_{1}\dots n_{l}\cap j)$ be the probability of observing the output $(n_{1},\dots,n_{l})$ and detecting $j$ photons in all $m$ modes, with $j\in\{c,\dots,n\}$. When postselecting on no photons being lost, we are computing
@@
Thm. 16, proven in Appendix 15, is strong evidence that techniques based on Richardson extrapolation offer no advantage over postselection. Our main technical contribution in proving Thm. 16 is to link determining the error of ZNE methods to computing the norm of the inverse of a Vandermonde matrix whose entries are determined by $\eta$, we then use existing results to upper bound this norm [20].
To strengthen our theoretical analysis, we also numerically compute the number of violations of the inequality $E_{extrap}\geq\frac{\epsilon_{max}}{\sqrt{(1-\eta_{0})^{n}}}$ and plot these in Figure 6. We took $\epsilon_{max}=0.01$, $\eta_{0}=0.01$, $\eta_{n-c}=0.95$, and $\eta_{i}$ for 1<i<nc1<i<n-c equally spaced. We varied the value of ncn-c between 3 and 14, with $c=\mathsf{ceil}(n/3)$ and $\mathsf{ceil}(.)$ is the ceiling function. For each value of ncn-c we performed 3000 runs, where at each run we took nc+1n-c+1 values of $\{\epsilon_{i}\}$ chosen uniformly randomly from $\epsilon_{max},\epsilon_{max}]$. As can be observed in Figure 6, the number of violations approaches zero with increasing $n$, confirming that extrapolation performs worse than post-selection after some value of $n$. A similar behaviour is observed for different values of $\epsilon_{max},\eta_{0}$ and $\eta_{n-c}$.
Figure 6: Number of violations of Thm. 16 inequality plotted versus ncn-c .
8 Discussion
In summary, we have presented a family of techniques, collectively referred to as recycling mitigation, for mitigating the effects of photon loss on the outputs of linear optical quantum circuits in the discrete variable setting. We provided analytical and numerical evidence that these techniques outperform postselection - currently the standard method for mitigating loss in linear optical circuits.
There are many possible directions of future work. For instance, carrying out of a rigorous analysis of the set-up where recycling mitigation is applied to linear optical computation with adaptive measurements and feedforward, such as in [13]. This would be interesting for a number of reasons, not least of which is that in this computational set-up, with a sufficient number of adaptive measurement steps, a quantum advantage for probability estimation is feasible [13]. As an estimate, if $n_{ad}$ is the number of photons consumed in the $r$ feed-forward steps in the absence of loss, and $\eta$ is the loss per mode, the overall probability that the $r$ feed-forward steps are executed correctly is $(1-\eta)^{n_{ad}}$. This introduces an unavoidable additional exponential sampling overhead to recycling mitigation. Nevertheless, it would be interesting benchmark the performance of recycling mitigation against only postselection using a set-up of this kind, especially in the possible regime where there is not sufficient feed-forward for universality, but where classical simulation methods are also inefficient [13].
It would also be interesting to investigate how the method might be combined with early fault-tolerance schemes [29] for optical quantum computing. Specifically, in investigating the resource trade-off between error mitigation and correction needed to achieve a certain quality of computational output.
Another interesting question, motivated by finding ways to eliminate the bias error in our introduced mitigation techniques is whether one can exactly represent an nn–photon probability as a sum of nkn-k–photon probabilities. This can be done for the case of minors of unitary matrices with real and positive entries, such as those used to solve graph problems in DVLOQC [43]. As an example, consider a Laplace expansion of the permanent of an $n\times n$ minor $U_{\mathbf{s},\mathbf{t}}:=(u_{ij})_{i,j\in\{1,\dots,n\}}$ of a linear optical unitary $U$, this reads
If $U_{\mathsf{s},\mathsf{t}}$ is composed of positive real entries, then $\mathsf{Per}(U_{\mathsf{s},\mathsf{t}})$ is proportional to the square root of a probability $\sqrt{p(s|t)}$ of a certain output of linear optical circuit $U$ with $n$ input photons [1]. In turn, each of the permanents of $n-1\times n-1$ matrices on the right hand side of the above equation is proportional to the square root of a probability of a certain output of a linear optical circuit with n1n-1 input photons. Thus, in the case of minors with positive real entries, there is an exact way to represent an nn–photon probability as a sum of n1n-1 photon probabilities. In this case as well, the process can be generalised, by successive Laplace expansions, to expressing nn–photon probabilities exactly as a sum of nkn-k–photon probabilities.
This idea of exactly representing nn–photon probabilities as sums of nkn-k–photon probabilities is interesting since, in the presence of photon loss, nkn-k–photon experiments in general produce more converged statistics than nn–photon experiments, for a comparable number of runs of the experiments in both cases. Furthermore, since the nn–photon probabilities are exactly expressible as a sum of nkn-k–photon probabilities, one can use the nkn-k–photon experiments to compute the nn–photon probabilities and indefinitely (because of lower statistical error) outperform postselection. It is an interesting question to determine whether similar results hold for broader classes of matrices.
Finally, it could be interesting to extend our techniques to the case where we allow collisions in the output. The construction of the recycled probabilities presented generalises straightforwardly to the collision regime, the main technical challenge is therefore to adapt our proofs to this regime.
Code availability
Open-source code to perform recycling mitigation may be found at the following Github location:
Acknowledgments
Appendix
9 Results on useful regimes of photon loss
9.1 Proof of Lemma 6
We now prove Lemma 6 from the main text:
Let k=nrk=n-r, there is a classical algorithm running in time $O(2^{r-1}r({n\choose n-r})^{2})$ which exactly computes $\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i})$.
Proof.
10 Statistical error bounds for probability estimators
10.1 Proof of Lemma 17